1. The frequency of a sound wave was measured to be 331 Hz. If you know that the speed of sound in air was 343 m/s, then determine the wavelength of the sound wave.

2. The wavelength of a sound wave was measured to be 1.47 m. If you know that the speed of sound in air was 343 m/s, then determine the frequency of the sound wave.

3. If the wavelength and frequency of a sound wave was measured to be 0.98 m and 340 Hz, then determine the speed of sound.

4. A 1.04 m length tube was used to measure the speed of sound in air. If the time between the generated and detected sound was measured to be 6.24 ms, then determine the speed of sound in air.

5. If the temperature inside the lab was 23oC, then determine the speed of sound of air in the lab. Remember to convert the temperature from Celsius to Kelvin. (k = 1.38×10-23 J/K, m = 4.80×10-26 kg, and γ= 1.4)
Answer #1

1.

speed=wavelength*frequency

Wavelength=speed/frequency

λ=v/f

λ=(343m/s)/(331Hz)

λ= 1.04 m

2.

f=v/λ = 343/1.47 = 233.3 Hz

3)

v=f λ = 0.98*340 = 333.2 m/s